OpenStudy (anonymous):

(2500)(8000) = 20 X 10^x x =??

OpenStudy (anonymous):

\[\ln(2500*8000)=\ln(20*10^{x})\]

OpenStudy (anonymous):

that's the first step

OpenStudy (anonymous):

no use of calculator @esamalaa

OpenStudy (anonymous):

are you want x as a natural no.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so you have to use a calculator. i think :)

OpenStudy (anonymous):

(2500∗8000)=(20∗10^x) 25 x 8 x 100000 = 200 x 100000 = 20 x 1000000 = 20 x 10^6 {x=6} ans ????

OpenStudy (anonymous):

yea that's right :) good job